OpenStudy (anonymous):

Let's let the equation y = 0.50(1.06)x represent the price of movie tickets for the years after 1950. You have filled in the ticket table for the prices for the years 0, 10, 20, 30, 40. Do these prices look reasonable?

OpenStudy (anonymous):

are you sure you have the expression correct? I can't see why is isn't written as 0.53, rather than 0.50 (1.06) - why aren't they multiplied?

OpenStudy (anonymous):

y=0.50(1.06)^x

OpenStudy (anonymous):

Ahhh, THAT makes more sense! in 1950, price = 0.50 (1.06)^0 = 0.5 = $0.05 in 1960, price = 0.50 (1.06)^10 = 0.8954... = $0.90 in 1970, price = 0.50 (1.06)^20 = 1.6035.... = $1.60 in 1980, price = 0.50 (1.06)^30 = 2.8717.... = $2.87 in 1990, price = 0.50 (1.06)^40 = 5.1428.... = $5.14 in 2000, price = 0.50 (1.06)^50 = 9.2100.... = $9.21 in 2010, price = 0.50 (1.06)^60 = 16.4938.... = $16.49 no bad, tickets here now range from around $9 to $15

OpenStudy (anonymous):

do u think these prices look resonable?